OpenStudy (anonymous):

You have been asked to design a "ballistic spring system" to measure the speed of bullets. A spring whose spring constant is is suspended from the ceiling. A block of mass hangs from the spring. A bullet of mass is fired vertically upward into the bottom of the block. The spring's maximum compression is measured.

OpenStudy (anonymous):

draw the diagram and write down wat u want to find

mathslover (mathslover):

mv = (M + m) v' 1/2(M+m)v'^2 = 1/2 kd^2 v' = d sqrt(k/(M+ m)) v = d*sqrt[k*(m+M)]/m v_B = 204.77 m/s

OpenStudy (anonymous):

write down the data what u want to find and the jey to this problem is transfer of kinetic energy of bullet to potential energy possesd by spring

OpenStudy (anonymous):

I dnt need to include data. I just need the general form. The above stated answer is wrong. The answer according to mastering physics is

OpenStudy (anonymous):

waqt is that expression for? vvelocity of bullet initially?

OpenStudy (anonymous):

ohh sorry I forgot to type the other half of this question. The other half is Find an expression for the bullet's speed . Express your answer in terms of the variables m,M,k,d and constant g .

OpenStudy (anonymous):

can u draw the diagram it will make things simpler for us

OpenStudy (anonymous):

why do you need diagram for? The diagram is not given in the problem and the question is pretty clear to understand.

OpenStudy (anonymous):

then wat is d like that in the expression mark the data in the diagram then use gravity concept to obtain energy when it hits the mass